OpenStudy (anonymous):

the width of a rectangle is 6 centimeters less than the width. the perimeter is 64 centimeters.find the width

OpenStudy (anonymous):

i just want to know if i am on the right path p=2L + 2W w=L - 6

OpenStudy (anonymous):

if you meant "The LENGTH of a rectangle is 6 centimeters less than the width", then yes you are on the right track!

OpenStudy (anonymous):

no the question is exactly how i had writen it so i think im doing it wrong. may you help please :)

OpenStudy (ryan123345):

Hey there! I see that you are new to OpenStudy. This is a great website on getting help with questions of any subject. It is a fun way of learning! If you have any additional questions involving the website feel free to ask the people with the purples A's next to their name. They are the ambassadors and will answer all your questions regarding the site. And also read the http://openstudy.com/code-of-conduct so you know the rules and regulations :) Have a great day and I hope you have fun!

OpenStudy (anonymous):

Uh. Double check that, because you used the word width twice.

OpenStudy (anonymous):

^

OpenStudy (anonymous):

alrighty one second brb

OpenStudy (anonymous):

the question actually is the width of a rectangle is 6 centimeters less than the length.the perimeter is 64 centimeters. find the width

OpenStudy (anonymous):

Then what you had was close, just switch the L and the W... W = L - 6

OpenStudy (anonymous):

okay let me try

OpenStudy (anonymous):

p=2L+2W W=L-6 P=2L=2(L-6) 64=2L+2L-12 64=4L-12 76=4L 76/4=4/4 L=19 i have to plug in now correct to find w

OpenStudy (anonymous):

Correct, keep going!

OpenStudy (anonymous):

64=2(19) + 2w 64=38+ 2w 64-38=38-38+2w 26=2w 26/2=2w/2 13=w

OpenStudy (anonymous):

Congrats! You did it mate.

OpenStudy (anonymous):

thank you